# Towards Unified Task Embeddings Across Multiple Models: Bridging the Gap for Prompt-Based Large Language Models and Beyond

### v1

@@ Abstract
Task embedding, a meta-learning technique that captures task-specific information, has become prevalent, especially in areas such as multi-task learning, model editing, and interpretability. However, it faces challenges with the emergence of prompt-guided Large Language Models (LLMs) operating in a gradient-free manner. Existing task embedding methods rely on fine-tuned, task-specific language models, which hinders the adaptability of task embeddings across diverse models, especially prompt-based LLMs. To unleash the power of task embedding in the era of LLMs, we propose a framework for unified task embeddings (FUTE), harmonizing task embeddings from various models, including smaller language models and LLMs with varied prompts, within a single vector space. Such uniformity enables the comparison and analysis of similarities amongst different models, extending the scope and utility of existing task embedding methods in addressing multi-model scenarios, whilst maintaining their performance to be comparable to architecture-specific methods.
@@ 1 Introduction
Figure 1: Overview of FUTE. Existing methods generate task embeddings by leveraging model parameters, confining to models of identical architecture and initialization. This limitation, depicted on the left, results in embeddings that reside in separate vector spaces depending on the model, making cross-model similarity comparisons impossible. Our unified framework introduces a framework by separating and independently computing task embeddings, enabling the incorporation of diverse models, ranging from different neural architectures to Large Language Models (LLMs) with distinct prompts, into a unified vector space.
@@
Task embeddings serve as a valuable meta-learning tool for various downstream tasks such as multi-task learning (Vu et al., 2020; Li et al., 2022; Zhou et al., 2022; Wu et al., 2023), model editing (Ilharco et al., 2023; Ortiz-Jimenez et al., 2023), and interpretability (Gueta et al., 2023). Current task embedding methods focus on encapsulating the task-specific information into a vector representation, which is derived from parameters of task-specific, fine-tuned model (Vu et al., 2020; Zhou et al., 2022; Ilharco et al., 2023). However, the existing paradigm’s reliance on the parameters of a single model restricts the comparison and adaptability of task embeddings across a range of models, especially hindering their applicability to prompt-guided, gradient-free LLMs.
@@
To resolve such hurdles, we introduce a model-agnostic Framework for Unified Task Embedding (FUTE), which enables the application of task embedding to diverse models including language models of various architectures to LLMs with different prompts, as illustrated in Figure 1. FUTE generates task embeddings by treating these models as “black boxes” and does not require any knowledge of the original models’ training dataset. Specifically, we decouple the concept of task embedding into two distinct forms: Dataset Task Embedding (DTE) and Model Task Embedding (MTE). DTE is designed to capture the task information inherent to a specific dataset, focusing on the dataset’s unique characteristics. On the other hand, MTE captures the model’s behaviour in accordance with the task performed, and it is independent of their training dataset. By decoupling MTE and DTE, our framework facilitates a more granular analysis of task characteristics, allowing for a deeper understanding of both the data and models employed. Moreover, the DTEs and MTEs of various models are computed using a single surrogate model, which is employed to substitute the role that the original model plays in existing methods. Such a uniform manner of computing DTEs and MTEs keeps the usage of existing methods unchanged. This ensures that the learned task embeddings are residing in the same embedding space, hence facilitating the comparison and analysis of similarities among them.
@@
Building upon aforementioned improvements, FUTE conceptualizes the combination of a prompt and a LLM as a model unit, allowing for the computation of task embeddings for LLMs that are guided to perform specific tasks through prompts. Experiments show that, while being more adaptable and covering a wider range of models, our proposed framework performs on par with existing model-specific embedding methods. Our main contributions are summarized as follows:
@@
  • Experiments show that FUTE, while being more versatile, retains a performance to be comparable to the existing model-specific methods.
@@ 2 Related Work
Task embedding, initially introduced as a task vector (Achille et al., 2019), as a meta-learning method for visual classification tasks. This concept utilizes the empirical Fisher information derived from model gradients to encode task-specific information into a vector representation, demonstrating that the model parameter can provide detailed task information. Expanding into NLP, Vu et al. (2020) applied Fisher information to capture task-specific details in fine-tuned language models, formalizing it as task embedding. Zhou et al. (2022) explored three Parameter-Efficient Fine-Tuning (PEFT) methods, Prompt Tuning (Liu et al., 2022b), Bias Tuning (Ben Zaken et al., 2022), and Low-Rank Tuning (Hu et al., 2022) to generate task embeddings. They discovered that PEFT parameters, being the only components adjusted during fine-tuning, inherently encapsulate task-specific information. Ilharco et al. (2023) introduced a method with arithmetic operations by subtracting the original parameters from its fine-tuned parameters to extract task vectors. Additional research efforts, such as (Lv et al., 2023; Huang et al., 2023a, b) have indirectly engaged with concepts akin to task embedding for multi-task learning, albeit without explicitly naming them as such. Recent investigations into the task embedding of LLMs by (Hendel et al., 2023) extracted task vectors from parameters and activations of LLMs. However, similar to previous approaches, this method is confined to a single model and is unable to generate task embeddings for datasets, thereby limiting its applicability. All the aforementioned methods predominantly focused on learning task embeddings for individual models, limiting their application in scenarios involving multiple models.
@@ 3.3 Model Task Embedding
For sourcing unsupervised text data $U$, we opt for the CrossFit dataset (Ye et al., 2021). CrossFit encompasses a dataset collection of 160 diverse NLP tasks, offering a rich variety of text styles and structures. This diversity stands in contrast to large, unsupervised text corpora such as Wikipedia, which predominantly feature formal and neutral tones, and consistent style. Such uniformity in text data can skew the performance of models towards, for example, neutral predictions in tasks such as sentiment analysis. This bias may require a much larger volume of data to develop a comprehensive understanding of the model’s capabilities. Given this context, the broad spectrum of text characteristics found in CrossFit makes it an ideal choice for $U$ in our study, especially when considering the efficiency of MTE learning. In our experiments, Eq. (5) with TuPaTE requires only 20 minutes when employing CrossFit as $U$. See Appendix A.4 for more details.
@@ 3.4 Dataset Task Embedding
Dataset Task Embedding (DTE) aims to encapsulate the task information of a dataset within an embedding. The process of learning DTE closely aligns with that of MTE utilizing $\Theta_{T}$. Specifically, as illustrated in Figure 2B, given a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, the computation of DTE $e_{D}$ is revised as follows:
@@ 4.1.1 Experimental Setup
Our approach to the experiment diverges from the original method by shifting the focus from selecting the most appropriate source dataset to identifying the most compatible models for a target dataset. This is achieved by matching the MTEs of candidate models with the DTE of the target dataset based on their embedding similarity. See Appendix A.4 for efficiency analysis. More experimental setup details are given below:
@@
Baselines
We compare FUTE with the following baselines: (1) DataSize (Vu et al., 2020) uses the size of dataset. (2) CurveGrad (Bingel and Søgaard, 2017; Vu et al., 2020) computes gradients for language model. (3) TextEmb (Vu et al., 2020) extracts text representations extracted from the language model. (4) TaskEmb (Vu et al., 2020) and TuPaTE (Zhou et al., 2022), as detailed in Section 3.2. See Appendix A.2 for more details.
@@ 4.2.1 Experimental Setup
In this section, we assess the effectiveness of FUTE for LLMs by focusing on the task of zero-shot prompt instruction selection. The objective is to identify the most suitable prompt instruction that enables an LLM to achieve optimal performance on a given dataset. We utilize FUTE to generate MTEs of LLMs associated with prompts using Eq. (5), and compute DTE for the target dataset using Eq. (6). We then compute the similarity between MTEs and DTEs for prompt selection. See Appendix A.8 for efficiency analysis.
@@
We select LLMs that represent a range of scales and architectures, including Llama2-13B, Llama2-70B (Touvron et al., 2023), and Mixtral 8x7B (Jiang et al., 2024). Mixtral 8x7B is a Sparse Mixture of Experts (SMoE) architecture, featuring 8 experts with each token processed by two experts.
@@ 4.2.2 Baselines
We compare FUTE against the following baselines: (1) MI (Sorensen et al., 2022) calcualtes the mutual information. (2) LocalE (Lu et al., 2022) and (3) GlobalE (Lu et al., 2022) compute the entropy for each entry and for entries across the dataset, respectively. ZPS (Liao et al., 2022) selects prompt by generating pseudo labels using ensembling with three variants: (4) ZPS-Log using the sum of log probabilities, (5) ZPS-Prob using the sum of probabilities, and (6) ZPS-Vote using majority vote. (7) Self-Select (Ramji and Kyimpopkin, 2023) uses a meta-prompt to query LLM selecting prompt. (8) SPELL (Gonen et al., 2023) computes the perplexity of prompts. See Appendix A.6 for more details.
@@ 5 Conclusion and Future Work
FUTE is designed to extend, rather than enhance, current task embedding learning methods. Consequently, it inherits potential limitations from the methodologies it builds on. As depicted in Figure 2, FUTE requires the use of an additional surrogate base model $\Theta_{T}$ and fine-tuning for learning task embedding, mirroring the paradigm of prior methods. The choice of this surrogate base model may affect the resultant task embeddings, as all embeddings are predicated on the model’s parameters to remain within the same vector space. Our current experiment depends on a surrogate T5 model to emulate the behaviors of the target model. However, given its smaller scale, T5 might not accurately reproduce the target model’s behavior. In the future, we could consider alternative surrogate base models. We acknowledge that any alteration to the surrogate base model requires the re-learning of task embeddings, which could be computationally intensive. However, we consider this issue intrinsic to the existing task embedding learning methods.
@@ A.1 Datasets in Language Model Experiment
A.2 Baselines in Language Model Experiment
All these baselines are unavailable in the multi-model scenario. Dataset-centric methods, like DataSize, inherently cannot differentiate between models trained on the same dataset. In the multi-model scenario, multiple models will be trained on the same dataset. Assigning the same features for different models, even if they are trained on the same dataset, is unreasonable. Model-specific approaches, CurveGrad, TextEmb, TaskEmb, and TuPaTE, all require a single model initialization, limiting their adaptability across diverse models. CurveGrad computes the gradients from the same model. TextEmb extracts the text representations from the same model. For example, similarly, text embedding derived from BERT-base, BERT-large, and GPT2 do not reside in the same vector space. However, our framework can be potentially adapted to extend text embedding methods as well. We leave it in the future work. TaskEmb and TuPaTE, as mentioned above, are also unavailable for multiple models.
A.3 Visualization of Language Model Task Embedding
@@
A.4 Efficiency Analysis on Language Model
A key feature of FUTE is the utilization of a consistent unsupervised dataset, denoted as $U$, for MTEs across all models, ensuring uniform computational resource requirements. The learning of Eq. (5) with TaskEmb requires approximately 50 minutes, while TuPaTE integration requires around 20 minutes, on a NVIDIA A100 GPU. In contrast, the runtime of original TaskEmb and TuPaTE depends on the dataset size. For instance, processing a small dataset like WNLI (Levesque et al., 2011) can be accomplished in approximately 1 hour, whereas handling a larger dataset such as MNLI (Williams et al., 2018) may extend beyond 2 days. The consistent use of $U$ reduces variability in computational demand across different models and also demonstrates robust computational efficiency.
A.5 Prompt Template
@@ 11 Given that {premise} Does it follow that hypothesis yes, no, or maybe?
A.6 Baselines in LLM Experiment
In LLM experiment, we compare FUTE against the following baselines: (1) MI (Sorensen et al., 2022) selects the prompt that maximizes the mutual information between the input and the LLM output. (2) LocalE (Lu et al., 2022) calculates the local entropy of the prediction probabilities for each entry. (3) GlobalE (Lu et al., 2022) computes the global entropy across all prediction categories in a dataset. ZPS (Liao et al., 2022) leverages an ensemble of prompts to generate pseudo labels and then predicts the performance of each prompt on these pseudo labels. Based on the ensembling technique, it has three variants: (4) ZPS-Log using the sum of log probabilities, (5) ZPS-Prob using the sum of probabilities, and (6) ZPS-Vote using majority vote. (7) Self-Select (Ramji and Kyimpopkin, 2023) introduces a meta-prompt that queries LLM for selecting the most effective prompt. (8) SPELL (Gonen et al., 2023) predicts prompt performance based on the perplexity of the prompt.
A.7 Visualization of LLM Task Embedding with Prompts
@@
A.8 Efficiency Analysis on LLMs
FUTE extends the existing task embedding method and utilizes embedding similarity to select prompts, leading to a distinctive computational complexity profile. The process requires evaluating a total of $k_{p}$ prompts across $k_{D}$ datasets. For identifying the optimal prompt, the computational complexity of FUTE is represented as $O((k_{p}+k_{D})c_{o})$, where $c_{o}$ represents the computation cost associated with the task embedding method. In contrast, the computational complexity for other baselines evaluated in this study, denoted as $c_{b}$, is expressed as $O(k_{p}k_{D}c_{b})$. This difference in complexity becomes particularly significant with increasing values of $k_{p}$ and $k_{D}$, where the product $k_{p}k_{D}$ will substantially exceed the sum $(k_{p}+k_{D})$. Thus, FUTE presents a potentially more efficient approach to prompt selection, especially as the number of prompts and datasets grows. However, it is crucial to note that the existing task embedding methods used in FUTE require fine-tuning, thereby $c_{o}>c_{b}$. With a relatively small number of prompts ($k_{p}$) and datasets ($k_{D}$), FUTE does not demonstrate superior running efficiency in comparison to baselines that require less computational effort per operation. Nonetheless, $c_{o}$ is more closely related to the inherent efficiency of the task embedding method itself, whereas FUTE primarily focuses on extending these methods.

### v2

@@ Abstract
Task embedding, a meta-learning technique that captures task-specific information, has gained popularity, especially in areas such as multi-task learning, model editing, and interpretability. However, it faces challenges with the emergence of prompt-guided Large Language Models (LLMs) operating in a gradient-free manner. Existing task embedding methods rely on fine-tuned, task-specific language models, which hinders the adaptability of task embeddings across diverse models, especially prompt-based LLMs. To hardness the potential of task embeddings in the era of LLMs, we propose a framework for unified task embeddings (FUTE), harmonizing task embeddings from various models, including smaller language models and LLMs with varied prompts, within a single vector space. Such uniformity enables comparison and analysis of similarities amongst different models, broadening the scope and utility of existing task embedding methods in multi-model scenarios, while maintaining their performance comparable to architecture-specific methods. [^1] Code is available at https://github.com/xnyuwg/fute.
@@ 1 Introduction
Figure 1: Overview of FUTE. Existing methods generate task embeddings by leveraging model parameters, confining to models of identical architecture and initialization. This limitation, depicted on the left, results in embeddings that reside in separate vector spaces depending on the model, making cross-model similarity comparisons impossible. Our unified framework introduces a method for independently computing task embeddings, enabling the incorporation of diverse models, ranging from different neural architectures to Large Language Models (LLMs) with distinct prompts, into a unified vector space.
@@
Task embeddings serve as a valuable meta-learning tool for various downstream tasks such as multi-task learning (Vu et al., 2020; Li et al., 2022; Zhou et al., 2022; Wu et al., 2023), model editing (Ilharco et al., 2023; Ortiz-Jimenez et al., 2023), and interpretability (Gueta et al., 2023). Current task embedding methods focus on encapsulating the task-specific information into a vector representation, which is derived from parameters of task-specific, fine-tuned models (Vu et al., 2020; Zhou et al., 2022; Ilharco et al., 2023). However, the existing paradigm’s reliance on the parameters of a single model restricts the comparison and adaptability of task embeddings across various models, especially hindering their applicability to prompt-guided, gradient-free Large Language Models (LLMs).
@@
To overcome these challenges, we propose a model-agnostic Framework for Unified Task Embedding (FUTE), which enables the application of task embedding across diverse models including language models of various architectures and LLMs with different prompts, as illustrated in Figure 1. FUTE generates task embeddings by treating these models as “black boxes” and does not require knowledge of the original models’ training datasets. Specifically, we decouple the concept of task embedding into two distinct forms: Dataset Task Embedding (DTE) and Model Task Embedding (MTE). DTE is designed to capture the task information specific to a dataset, focusing on its unique characteristics. In contrast, MTE captures the model’s behavior according to the task performed, independent of its training dataset. By decoupling MTE and DTE, our framework facilitates a more granular analysis of task characteristics, providing insights into both the data and models employed. Moreover, the DTEs and MTEs of various models are computed using a single surrogate model, which replaces the role of the original model in existing methods. This consistent approach to computing DTEs and MTEs ensures that the learned task embeddings reside in the same embedding space, hence facilitating the comparison and analysis of similarities among them.
@@
Building upon the aforementioned improvements, FUTE conceptualizes the combination of a prompt and a LLM as a single model unit, enabling the computation of task embeddings for LLMs guided to perform specific tasks through prompts. Experimental results show that, while being more adaptable and covering a wider range of models, our proposed framework performs on par with existing model-specific embedding methods. FUTE primarily focuses on extending task embedding methods to multi-model scenarios rather than directly improving the embedding methods themselves. Although the performance of FUTE relies on the employed task embedding methods, this aspect represents a strength, allowing FUTE to leverage future advancements in task embedding techniques without significant modifications. Our main contributions are summarized as follows:
@@
  • Experimental results show that FUTE, while being more versatile, retains a performance to be comparable to existing model-specific methods.
@@ 2 Related Work
Task embedding, initially introduced as a task vector (Achille et al., 2019), as a meta-learning method for visual classification tasks. This concept utilizes the empirical Fisher information derived from model gradients to encode task-specific information into a vector representation, demonstrating that the model parameter can provide detailed task information. Expanding into NLP, Vu et al. (2020) applied Fisher information to capture task-specific details in fine-tuned language models, formalizing it as task embedding. Zhou et al. (2022) explored three Parameter-Efficient Fine-Tuning (PEFT) methods, Prompt Tuning (Liu et al., 2022b), Bias Tuning (Ben Zaken et al., 2022), and Low-Rank Tuning (Hu et al., 2022) to generate task embeddings. They discovered that PEFT parameters, being the only components adjusted during fine-tuning, inherently encapsulate task-specific information. Ilharco et al. (2023) introduced a method with arithmetic operations by subtracting the original parameters from its fine-tuned parameters to extract task vectors. Studies by (Alvarez-Melis and Fusi, 2020; Tan et al., 2021) proposed direct computation of task similarity via Wasserstein Distance, bypassing the need for embedding similarity measures. Additional research efforts, such as (Lv et al., 2023; Huang et al., 2023a, b) have indirectly engaged with concepts akin to task embedding for multi-task learning, albeit without explicitly naming them as such. Recent investigations into the task embedding of LLMs by (Hendel et al., 2023) extracted task vectors from parameters and activations of LLMs. However, similar to previous approaches, this method is confined to a single model and is unable to generate task embeddings for datasets, thereby limiting its applicability. All the aforementioned methods predominantly focused on learning task embeddings for individual models, limiting their application in scenarios involving multiple models.
@@ 3.3 Model Task Embedding
For sourcing unsupervised text data $U$, we opt for the CrossFit dataset (Ye et al., 2021). CrossFit encompasses a dataset collection of 160 diverse NLP tasks, offering a rich variety of text styles and structures. This diversity stands in contrast to large, unsupervised text corpora such as Wikipedia, which predominantly feature formal and neutral tones, and consistent style. Such uniformity in text data can skew the performance of models towards, for example, neutral predictions in tasks such as sentiment analysis. This bias may require a much larger volume of data to develop a comprehensive understanding of the model’s capabilities. Given this context, the broad spectrum of text characteristics found in CrossFit makes it an ideal choice for $U$ in our study, especially when considering the efficiency of MTE learning.
@@ 3.4 Dataset Task Embedding
Dataset Task Embedding (DTE) aims to encapsulate the task information of a dataset within an embedding. In this paper, DTE is conceptualized similarly to the task embedding methods presented in prior studies. The process of learning DTE closely aligns with that of MTE utilizing $\Theta_{T}$. Specifically, as illustrated in Figure 2B, given a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, the computation of DTE $e_{D}$ is revised as follows:
@@
This reformulation facilitates a more flexible and adaptable approach for generating task embeddings, by leveraging $\Theta_{T}$ without being bounded by the constraints of architectural and initialization uniformity imposed by $\Theta_{M}$.
@@ 4.1.1 Experimental Setup
Our approach to the experiment diverges from the original method by shifting the focus from selecting the most appropriate source dataset to identifying the most compatible models for a target dataset. This is achieved by matching the MTEs of candidate models with the DTE of the target dataset based on their embedding similarity. More experimental setup details are given below:
@@
4.1.2 Baselines
All these baselines are unavailable in the multi-model scenario. Dataset-centric methods, like DataSize, inherently cannot differentiate between models trained on the same dataset. Model-specific approaches, CurveGrad, TextEmb, TaskEmb, and TuPaTE, all require a single model initialization, limiting their adaptability across diverse models. CurveGrad computes the gradients from the same model. TextEmb extracts the text representations from the same model. For example, similarly, text embedding derived from BERT-base, BERT-large, and GPT2 do not reside in the same vector space. However, FUTE can be potentially adapted to extend text embedding methods as well. TaskEmb and TuPaTE, as mentioned above, are also unavailable for multiple models.
@@ 4.2.1 Experimental Setup
In this section, we assess the effectiveness of FUTE for LLMs by focusing on the task of zero-shot prompt instruction selection. The objective is to identify the most suitable prompt instruction that enables an LLM to achieve optimal performance on a given dataset. We utilize FUTE to generate MTEs of LLMs associated with prompts using Eq. (5), and compute DTE for the target dataset using Eq. (6). We then compute the similarity between MTEs and DTEs for prompt selection.
@@
We employ Llama2-13B (Touvron et al., 2023) in 8 bit, Llama2-70B (Touvron et al., 2023) in 4 bit, and Mixtral 8x7B (Jiang et al., 2024) in 8 bit. Mixtral 8x7B is a Sparse Mixture of Experts (SMoE) architecture, featuring 8 experts with each token processed by two experts.
@@ 4.2.2 Baselines
We compare FUTE against the following baselines: (1) MI (Sorensen et al., 2022) selects the prompt that maximizes the mutual information between the input and the LLM output. (2) LocalE (Lu et al., 2022) calculates the local entropy of the prediction probabilities for each entry. (3) GlobalE (Lu et al., 2022) computes the global entropy across all prediction categories in a dataset. ZPS (Liao et al., 2022) leverages an ensemble of prompts to generate pseudo labels and then predicts the performance of each prompt on these pseudo labels. Based on the ensembling technique, it has three variants: (4) ZPS-Log using the sum of log probabilities, (5) ZPS-Prob using the sum of probabilities, and (6) ZPS-Vote using majority vote. (7) Self-Select (Ramji and Kyimpopkin, 2023) introduces a meta-prompt that queries LLM for selecting the most effective prompt. (8) SPELL (Gonen et al., 2023) predicts prompt performance based on the perplexity of the prompt.
@@ 4.2.4 Visualization
4.2.5 Efficiency Analysis
FUTE extends the existing task embedding method and utilizes embedding similarity to select prompts, leading to a distinctive computational complexity profile. The process requires evaluating a total of $k_{p}$ prompts across $k_{D}$ datasets. The computational complexity of FUTE is represented as $O((k_{p}+k_{D})c_{o})$, where $c_{o}$ represents the computation cost associated with the task embedding method. In contrast, the computational complexity for other baselines evaluated in this study is expressed as $O(k_{p}k_{D}c_{b})$, where $c_{o}$ represents the computation cost associated with baselines. The difference in complexity becomes significant with increasing values of $k_{p}$ and $k_{D}$, where the product $k_{p}k_{D}$ will exceed the sum $(k_{p}+k_{D})$. Thus, FUTE presents a potentially more efficient approach to prompt selection, especially as the number of prompts and datasets grows. However, the existing task embedding methods used in FUTE require fine-tuning, thereby $c_{o}>c_{b}$. With a relatively small number of prompts ($k_{p}$) and datasets ($k_{D}$), FUTE does not demonstrate superior running efficiency in comparison to baselines. Nonetheless, $c_{o}$ is more closely related to the inherent efficiency of the task embedding method itself, whereas FUTE focuses on extending these methods.
@@ 5 Conclusion and Future Work
FUTE is designed to extend, rather than enhance, current task embedding learning methods. Consequently, it inherits potential limitations from the methodologies it builds on. However, we argue that this is more of a feature of FUTE rather than a limitation. By not altering current procedures, our framework leverages the strengths of existing task embedding methods, allowing for seamless integration of future advancements in the field. This adaptability means that as stronger task embedding methods emerge in the future, our framework will also exhibit improved performance without requiring modifications. We also acknowledge that any alteration to the surrogate base model requires the re-learning of task embeddings, which could be computationally intensive. However, we consider this issue intrinsic to the existing task embedding learning methods. If future developments in task embedding learning methods eliminate dependency on the base model, FUTE can be adjusted by simply omitting the surrogate base model as well.
@@ A.1 Datasets in Language Model Experiment
A.2 Visualization of Language Model Task Embedding
@@
A.3 Prompt Template
@@ 11 Given that {premise} Does it follow that hypothesis yes, no, or maybe?
A.4 Visualization of LLM Task Embedding with Prompts